Find the slope and the y-intercept of the line.
4x=5y+15

Answer by nabla(475)

Solve for y.
(4x-15)/5=y
(4/5)x-3=y

The slope of a line is m in y=mx+b. b is the y-intercept [b].
Thus, we get m=4/5, y-intercept=b=-3.